# A systematic review of metaheuristics-based and machine learning-driven intrusion detection systems in IoT

### v3

@@ 5.1.1 Population-based Metaheuristics and ML
Population-based Hybrid Metaheuristics. Numerous works have employed more than one population-based metaheuristics to improve the performance of IoT-IDSs. In most cases, such hybridization is utilized to select an optimal set of features. Regarding traditional machine learning classifiers, KNN [55_abu2021iot, 84_mahanipour2024enhancing, 1_kareem2022effective, 3_alkanhel2023network] and RF [13_krishna2021attack, 38_gaber2023industrial, 41_ethala2022hybrid, 33_phalguna2021hybrid, 44_keserwani2021smart] are widely used in these systems. Specifically, SSA+ALO [55_abu2021iot], quantum-driven binary ABC+GA [84_mahanipour2024enhancing], gorilla troops optimizer (GTO)+birds swarm algorithm (BSA) [1_kareem2022effective], and GWO+dipper throated optimization (DTO) [3_alkanhel2023network, 27_alkanhel2023hybrid] are integrated in these KNN-based detection systems. Notably, all these IDSs provide substantial performance in terms of accuracy, precision, recall, and other related metrics. On the other hand, RF-based IDSs utilize LOA+FOA [13_krishna2021attack], PSO+bat algorithm (BA) [38_gaber2023industrial], spider monkey algorithm+hierarchical PSO [41_ethala2022hybrid], and PSO+GWO [33_phalguna2021hybrid, 44_keserwani2021smart] for feature selection purpose. These systems also demonstrate significant efficiency in classifying various attacks. In addition to these, hunger game search (HGS) and remora optimization algorithm (ROA) in [31_kumar2022intellectual], and GA and GWO are hybridized in [30_davahli2020hybridizing] for selecting near-best features from AWID dataset for SVM classifier. Both of these combinations show notable performances with 99.1% accuracy and negligible FPR. Turning to deep learning-oriented IoT-IDSs, three works are found that applied a Look Ahead Artificial Neural Network (LAANN), recurrent neural networks (RNNs), and a deep learning-based hybrid neural network (DLHCNN) respectively. Particularly, sea turtle foraging algorithm (STFA)+explorated PSO (EXPSO) [54_jeyaselvi2023highly], Harris hawk optimization (HHO)+fractional derivative mutation (FDM) [7_sanju2023enhancing], and Chicken Swarm Optimization (ChSO)+GA [67_gupta2022hybrid] are utilized in these systems to select optimal feature set. However, they fail to provide remarkable performance with 95%98% considering all related metrics.
@@
Apart from feature selection-based IDSs, a few research studies focus on tuning the parameters or hyperparameters of classification models. Khafaga et al. [4_khafaga2023voting] propose an innovative whale optimization (WOA) regulated by DTO to optimize parameters of KNN, RF, and NN. Experimental evaluation using the RPL-NIDS17 dataset results in 99% AUC and 95.1% accuracy. In [26_sagu2023design], SAEHO and SU-CMO are also proposed for adjusting the parameters of the two hybrid classifiers, particularly CNN+DBN and Bi-LSTM+GRU. The integration of SAEHO and hybrid classifier offers better accuracy than that of SU-CMO and hybrid classifier (91.6% vs 84.8% ). Bahaa et al. [29_bahaa2022novel] integrate adaptive PSO and WOA for hyperparameter optimization in their CNN-based detection system. However, the system achieves only 94.54% accuracy and 0.9 JSC. However, a small number of papers focus on employing hybrid population-based optimizers for accomplishing multiple purposes at a time [45_karthikeyan2024firefly, 40_alruwaili2023red, 86_aburasain2024enhanced]. One of the notable systems is introduced by Karthikeyan et al. [45_karthikeyan2024firefly], where GWO is leveraged to optimize parameters and FOA to choose the most suitable features for the SVM classifier in the IoT-WSN environment. The separate use of these two metaheuristics crucially influences the system’s accuracy (99.29%). However, other such existing IoT-IDSs drastically fails, especially red kite optimization algorithm (RKOA)+Levy flight chaotic WOA with EL (LSTM, BiLSTM, and Bi-GRU) [40_alruwaili2023red] and black widow optimization (BWO)+BES with hybrid deep learning (HDL) [86_aburasain2024enhanced].
5.1.2 Iterative-based Metaheuristics and ML
@@
Physics and Math-based Regarding math-inspired optimization techniques, the arithmetic optimization algorithm (AOA) is widely employed. In [19_fraihat2023intrusion], AOA is utilized to select optimal feature sets for random forest and extra trees classifiers. Experimental evaluations on four public datasets reveal that the IDS produces a much less false positive rate (0.002%) for the tests conducted using the NF-ToN-IoT-v2 dataset. Makhadmeh et al. [102_makhadmeh2024intrusion] also apply AOA for executing the same purpose using different classifiers, KNN. Interestingly, the accuracy of these two systems is almost identical (around 99.9%). Though AOA and quantum-driven PSO (QPSO) are utilized in [11_malibari2022novel] for different purposes, the deep wavelet neural network (DWNN)-based detection system does not secure substantial accuracy (98.21%). Turning to the physics-based IoT-IDSs, Atom Search Optimization (ASO) and Equilibrium Optimization (EO) techniques are utilized in [110_maazalahi2024k] to select optimal features prior to applying K-means clustering. Importantly, this hybridization demonstrates remarkable performance on NSL-KDD, UNSW-NB15, and KDD-CUP99 datasets. In another paper [113_li2024cooperative], a black hole optimization technique is employed to select near-optimal features from the UNSW-NB15 and NSL-KDD datasets. Two CNNs are utilized in parallel to identify intrusion and result in around 97.5%99.89% efficiency in terms of common performance metrics.
@@
Human-inspired. Considering the integration of metaheuristics adopted from human behaviors and their decision-making process, [20_alghamdi2022hybrid, 52_pan2021lightweight, 18_fatani2023enhancing, 115_otoum2021ids] IoT-IDSs are the most notable ones. Specifically, the political optimizer (PO) utilized for parameter tuning in a cascade forward neural network (CFNN) model, compact SCA (CSCA) for adjusting the parameters of the KNN classifier, and modified growth optimizer (GO) for selecting near-best features before training by CNN are significantly performed well with an accuracy of 99.86%, 98.27%99.327%, and 99.941%), respectively. However, these groups of optimizers are not always effective, especially for intrusion detection in IoT environment [28_forestiero2021metaheuristic, 71_basheri2023quantum, 105_10541183].
@@
Search-based Though IoT-IDSs of this category tend to bring significant performances, they are too rare in the literature. Only one of them is found [58_nazir2023novel], the authors utilize tabu search with the idea of cellular automata to succeed in the features selection task. Consequently, it results in accuracy, precision, and FPR of 99.5%, 97.92%, and 0.004%, respectively while classified using an ensemble learning method.
5.1.3 Hybrid Metaheuristics and ML
@@
In the context of IoT intrusion detection systems, existing hybrid optimization algorithms can be categorized into 6 small groups: swarm+physics-based [32_habib2020modified], nature+physics-inspired [14_dey2023metaheuristic, 21_alweshah2022intrusion], nature+math-inspired [24_rahmani2024improvement, 36_aljebreen2023binary], nature+human-based [81_hanafi2024intrusion, 83_singh2024intrusion], swarm+search-based [35_baniasadi2022novel], and nature+search-inspired [62_alghanam2023improved, 63_anuradha2022intrusion, 77_om2021harmony]. Though utilizing physics-based techniques with nature-inspired ones (binary gravitational search (BGSA)+GWO [14_dey2023metaheuristic], simulated annealing (SA)+shuffled shepherd optimization (SSO) [21_alweshah2022intrusion]) prove to be effective, integrating with swarm-based optimizers (multi-object PSO+Lévy flight [32_habib2020modified]) turns out to be inefficient. Regarding the third group, Rahmani et al. [24_rahmani2024improvement] employ grasshopper optimization (GAO) and AOA for parameters and hyperparameters tuning in a random neural network (RdNN), which generates an IDS having 99.56% precision and 99.37% detection rate. In another work [36_aljebreen2023binary], a binary chimp optimization algorithm (BCOA) is integrated with the sine cosine algorithm (SCA) for securing the IoT-WSN network. Although the IDS generates high accuracy and specificity (99.63% and 99.67%, respectively), the f1-score is not convincing (94.52%). The well-known human-based metaheuristic object-based learning (OBL) is hybridized with Harris Hawk Optimization (HHO) and Golden Jackal Optimization Algorithm (GJOA), respectively in [83_singh2024intrusion] and [81_hanafi2024intrusion] to select best possible features. These DT-based and LSTM-driven IoT-IDSs provide remarkable performances with 99.65%100% and 98.93% accuracy, respectively. Baniasadi et al. [35_baniasadi2022novel] adjust parameters of deep CNN (DCNN) utilizing neighborhood search (NS)-based PSO. They get a negligible amount of mean square error (0.00053%) with 98.86% accuracy and 95.32% specificity on the UNSW-NB15 and Bot-IoT datasets. Turning to the last group of this category, it can be concluded that selecting features using a combination of nature-inspired and search-based metaheuristics is not effective according to the results provided in [62_alghanam2023improved, 63_anuradha2022intrusion, 77_om2021harmony]. It is worth mentioning that these IDSs use EL classifier, variational autoencoder (VAE), and deep RL, respectively.
@@
Table 5 shows the list of works in the existing literature that leverage different metaheuristics-based techniques and ML algorithms. According to the table, most systems leverage nature-inspired optimization techniques to increase the efficiency of the classifiers. Besides, swarm-based, population-based hybrid, and hybrid metaheuristics are utilized significantly. Moreover, a thorough analysis of the IoT-IDSs based on different performance metrics, metaheuristics, their applications, ML algorithms, classification types, and datasets are presented in Appendix A (Table LABEL:tab:performance_appendix).
@@
| Swarm-based | [5_saif2022hiids, 23_alweshah2023intrusion, 48_sarwar2022design, 111_dakic2024intrusion] | [39_vanitha2023improved, 57_prajisha2022efficient, 111_dakic2024intrusion] | [9_vaiyapuri2023metaheuristics, 34_stankovic2022feature, 50_jovanovic2022feature, 53_qaddoura2024evolving] | [22_anusha2022intrusion, 69_kannan2020intrusion] | [17_elmasry2020evolving, 90_biju2024evaluated] | [17_elmasry2020evolving, 65_anushiya2023new, 71_basheri2023quantum, 74_saheed2023novel, 94_aljehane2024golden, 101_deore2022hybrid] | |
| Nature-inspired | [25_ghasemi2024new, 37_lv2023binary, 59_alamiedy2020anomaly, 68_sandhya2021enhancing, 87_harbi2024bio, 89_elsedimy2024novel, 92_sharmamulti] | [8_savanovic2023intrusion, 46_jayalatchumy2024improved, 56_xu2023iot, 60_gangula2023intrusion, 66_gadekallu2023moth, 78_gopalakrishnan2022new, 85_vijayan2024original, 91_saheed2024voting] | [6_li2022improving, 51_geetha2024cvs, 53_qaddoura2024evolving, 70_bathula2022designing, 79_mohy2023whale, 88_rabie2024novel, 104_10317882] | [63_anuradha2022intrusion, 64_om2022effective, 93_cherian2024iot, 98_moizuddin2022bio, 106_escorcia2023sea] | [2_rm2020effective, 16_dahou2022intrusion, 42_asgharzadeh2023anomaly, 43_jayasankar2024intrusion, 47_chander2023metaheuristic, 49_fatani2021advanced, 72_ramana2022wogru, 73_abd2023intrusion, 75_jothi2023wils, 80_bajpai2024hybrid, 82_kethineni2024intrusion, 94_aljehane2024golden, 95_bella2024intrusion, 96_vadakkethil2024mayfly, 99_kaviarasan2023network, 100_pingale2023remora, 107_10.1016/j.comcom.2023.03.005, 114_alamro2023modeling] | ||
| EA-based | [5_saif2022hiids, 12_dey2023hybrid] | [104_10317882, 112_latif2024dtl] | [65_anushiya2023new, 113_li2024cooperative] | ||||
| Population-based hybrid | [1_kareem2022effective, 3_alkanhel2023network, 4_khafaga2023voting, 13_krishna2021attack, 27_alkanhel2023hybrid, 30_davahli2020hybridizing, 31_kumar2022intellectual, 33_phalguna2021hybrid, 38_gaber2023industrial, 41_ethala2022hybrid, 44_keserwani2021smart, 45_karthikeyan2024firefly, 55_abu2021iot, 61_davahli2020lightweight, 84_mahanipour2024enhancing] | [4_khafaga2023voting, 54_jeyaselvi2023highly] | [26_sagu2023design] | [7_sanju2023enhancing, 26_sagu2023design, 29_bahaa2022novel, 40_alruwaili2023red, 67_gupta2022hybrid, 76_shahapure2021water, 86_aburasain2024enhanced, 114_alamro2023modeling] | |||
| Phy/Math-based | [19_fraihat2023intrusion, 102_makhadmeh2024intrusion] | [110_maazalahi2024k] | [11_malibari2022novel] | ||||
| Human-inspired | [52_pan2021lightweight, 97_10.1007/s11276-023-03435-0] | [20_alghamdi2022hybrid, 53_qaddoura2024evolving] | [105_10541183] | [18_fatani2023enhancing, 28_forestiero2021metaheuristic, 71_basheri2023quantum, 105_10541183] | |||
| Search-based | [58_nazir2023novel] | ||||||
| Hybrid | [21_alweshah2022intrusion, 32_habib2020modified, 83_singh2024intrusion] | [14_dey2023metaheuristic, 36_aljebreen2023binary, 62_alghanam2023improved] | [24_rahmani2024improvement] | [63_anuradha2022intrusion] | [77_om2021harmony] | [15_fatani2021iot, 35_baniasadi2022novel, 81_hanafi2024intrusion] | |
@@
We find that most of the existing IoT-IDSs solely utilize different metaheuristics algorithms for selecting an optimal set of features. Besides, some of the systems employ them distinctly for optimizing the parameters and hyperparameters of the machine learning models. However, a few works focus on hybrid applications by leveraging feature selection and parameter (FS-PT) or hyperparameter (FS-HPT) tuning in the same detection system. Figure 6 illustrates the percentage of different applications leveraged by the IDSs, and Table 6 shows the list of corresponding existing works.
@@
| FS | [1_kareem2022effective, 2_rm2020effective, 3_alkanhel2023network, 5_saif2022hiids, 6_li2022improving, 7_sanju2023enhancing, 12_dey2023hybrid, 13_krishna2021attack, 14_dey2023metaheuristic, 15_fatani2021iot, 16_dahou2022intrusion, 18_fatani2023enhancing, 19_fraihat2023intrusion, 21_alweshah2022intrusion, 22_anusha2022intrusion, 23_alweshah2023intrusion, 27_alkanhel2023hybrid, 30_davahli2020hybridizing, 31_kumar2022intellectual, 32_habib2020modified, 33_phalguna2021hybrid, 34_stankovic2022feature, 37_lv2023binary, 38_gaber2023industrial, 39_vanitha2023improved, 41_ethala2022hybrid, 42_asgharzadeh2023anomaly, 43_jayasankar2024intrusion, 44_keserwani2021smart, 46_jayalatchumy2024improved, 48_sarwar2022design, 49_fatani2021advanced, 50_jovanovic2022feature, 51_geetha2024cvs, 54_jeyaselvi2023highly, 55_abu2021iot, 56_xu2023iot, 57_prajisha2022efficient, 58_nazir2023novel, 59_alamiedy2020anomaly, 60_gangula2023intrusion, 61_davahli2020lightweight, 62_alghanam2023improved, 66_gadekallu2023moth, 67_gupta2022hybrid, 68_sandhya2021enhancing, 69_kannan2020intrusion, 73_abd2023intrusion, 76_shahapure2021water, 77_om2021harmony, 79_mohy2023whale, 81_hanafi2024intrusion, 83_singh2024intrusion, 84_mahanipour2024enhancing, 87_harbi2024bio, 88_rabie2024novel, 92_sharmamulti, 95_bella2024intrusion, 96_vadakkethil2024mayfly, 97_10.1007/s11276-023-03435-0, 98_moizuddin2022bio, 102_makhadmeh2024intrusion, 104_10317882, 105_10541183, 107_10.1016/j.comcom.2023.03.005, 110_maazalahi2024k, 113_li2024cooperative] |
| PT | [4_khafaga2023voting, 20_alghamdi2022hybrid, 24_rahmani2024improvement, 26_sagu2023design, 35_baniasadi2022novel, 53_qaddoura2024evolving, 64_om2022effective, 75_jothi2023wils, 82_kethineni2024intrusion, 90_biju2024evaluated, 93_cherian2024iot, 99_kaviarasan2023network, 100_pingale2023remora, 101_deore2022hybrid] |
| HPT | [8_savanovic2023intrusion, 24_rahmani2024improvement, 28_forestiero2021metaheuristic, 29_bahaa2022novel, 53_qaddoura2024evolving, 70_bathula2022designing, 71_basheri2023quantum, 72_ramana2022wogru, 80_bajpai2024hybrid, 85_vijayan2024original, 89_elsedimy2024novel, 91_saheed2024voting, 100_pingale2023remora, 111_dakic2024intrusion, 112_latif2024dtl] |
| FS and PT | [9_vaiyapuri2023metaheuristics, 11_malibari2022novel, 36_aljebreen2023binary, 45_karthikeyan2024firefly, 47_chander2023metaheuristic, 63_anuradha2022intrusion, 114_alamro2023modeling] |
| FS and HPT | [17_elmasry2020evolving, 25_ghasemi2024new, 36_aljebreen2023binary, 40_alruwaili2023red, 52_pan2021lightweight, 65_anushiya2023new, 74_saheed2023novel, 78_gopalakrishnan2022new, 86_aburasain2024enhanced, 94_aljehane2024golden, 106_escorcia2023sea] |
5.2 RQ2: What are the most commonly used datasets and evaluation metrics for IoT-IDS assessment?

### v4

@@ V-A1 Population-based Metaheuristics and ML
Population-based Hybrid Metaheuristics. Numerous works have employed more than one population-based metaheuristics to improve the performance of IoT-IDSs. In most cases, such hybridization is utilized to select an optimal set of features. Regarding traditional machine learning classifiers, KNN [2, 114, 91, 13] and RF [103, 65, 57, 132, 96] are widely used in these systems. Specifically, SSA+ALO [2], quantum-driven binary ABC+GA [114], gorilla troops optimizer (GTO)+birds swarm algorithm (BSA) [91], and GWO+dipper throated optimization (DTO) [13, 14] are integrated in these KNN-based detection systems. Notably, all these IDSs provide substantial performance in terms of accuracy, precision, recall, and other related metrics. On the other hand, RF-based IDSs utilize LOA+FOA [103], PSO+bat algorithm (BA) [65], spider monkey algorithm+hierarchical PSO [57], and PSO+GWO [132, 96] for feature selection purpose. These systems also demonstrate significant efficiency in classifying various attacks. In addition to these, hunger game search (HGS) and remora optimization algorithm (ROA) in [104], and GA and GWO are hybridized in [42] for selecting near-best features from AWID dataset for SVM classifier. Both of these combinations show notable performances with 99.1% accuracy and negligible FPR. Turning to deep learning-oriented IoT-IDSs, three works are found that applied a Look Ahead Artificial Neural Network (LAANN), recurrent neural networks (RNNs), and a deep learning-based hybrid neural network (DLHCNN) respectively. Particularly, sea turtle foraging algorithm (STFA)+explorated PSO (EXPSO) [87], Harris hawk optimization (HHO)+fractional derivative mutation (FDM) [150], and Chicken Swarm Optimization (ChSO)+GA [74] are utilized in these systems to select optimal feature set. However, they fail to provide remarkable performance with 95%98% considering all related metrics.
@@
Apart from feature selection-based IDSs, a few research studies focus on tuning the parameters or hyperparameters of classification models. Khafaga et al. [98] propose an innovative whale optimization (WOA) regulated by DTO to optimize parameters of KNN, RF, and NN. Experimental evaluation using the RPL-NIDS17 dataset results in 99% AUC and 95.1% accuracy. In [144], SAEHO and SU-CMO are also proposed for adjusting the parameters of the two hybrid classifiers, particularly CNN+DBN and Bi-LSTM+GRU. The integration of SAEHO and hybrid classifier offers better accuracy than that of SU-CMO and hybrid classifier (91.6% vs 84.8% ). Bahaa et al. [26] integrate adaptive PSO and WOA for hyperparameter optimization in their CNN-based detection system. However, the system achieves only 94.54% accuracy and 0.9 JSC. However, a small number of papers focus on employing hybrid population-based optimizers for accomplishing multiple purposes at a time [92, 16, 5]. One of the notable systems is introduced by Karthikeyan et al. [92], where GWO is leveraged to optimize parameters and FOA to choose the most suitable features for the SVM classifier in the IoT-WSN environment. The separate use of these two metaheuristics crucially influences the system’s accuracy (99.29%). However, other such existing IoT-IDSs drastically fails, especially red kite optimization algorithm (RKOA)+Levy flight chaotic WOA with EL (LSTM, BiLSTM, and Bi-GRU) [16] and black widow optimization (BWO)+BES with hybrid deep learning (HDL) [5].
V-A2 Iterative-based Metaheuristics and ML
@@
Physics and Math-based Regarding math-inspired optimization techniques, the arithmetic optimization algorithm (AOA) is widely employed. In [63], AOA is utilized to select optimal feature sets for random forest and extra trees classifiers. Experimental evaluations on four public datasets reveal that the IDS produces a much less false positive rate (0.002%) for the tests conducted using the NF-ToN-IoT-v2 dataset. Makhadmeh et al. [115] also apply AOA for executing the same purpose using different classifiers, KNN. Interestingly, the accuracy of these two systems is almost identical (around 99.9%). Though AOA and quantum-driven PSO (QPSO) are utilized in [116] for different purposes, the deep wavelet neural network (DWNN)-based detection system does not secure substantial accuracy (98.21%). Turning to the physics-based IoT-IDSs, Atom Search Optimization (ASO) and Equilibrium Optimization (EO) techniques are utilized in [113] to select optimal features prior to applying K-means clustering. Importantly, this hybridization demonstrates remarkable performance on NSL-KDD, UNSW-NB15, and KDD-CUP99 datasets. In another paper [108], a black hole optimization technique is employed to select near-optimal features from the UNSW-NB15 and NSL-KDD datasets. Two CNNs are utilized in parallel to identify intrusion and result in around 97.5%99.89% efficiency in terms of common performance metrics.
@@
Human-inspired. Considering the integration of metaheuristics adopted from human behaviors and their decision-making process, [9, 131, 59, 130] IoT-IDSs are the most notable ones. Specifically, the political optimizer (PO) utilized for parameter tuning in a cascade forward neural network (CFNN) model, compact SCA (CSCA) for adjusting the parameters of the KNN classifier, and modified growth optimizer (GO) for selecting near-best features before training by CNN are significantly performed well with an accuracy of 99.86%, 98.27%99.327%, and 99.941%), respectively. However, these groups of optimizers are not always effective, especially for intrusion detection in IoT environment [62, 29, 36].
@@
Search-based Though IoT-IDSs of this category tend to bring significant performances, they are too rare in the literature. Only one of them is found [127], the authors utilize tabu search with the idea of cellular automata to succeed in the features selection task. Consequently, it results in accuracy, precision, and FPR of 99.5%, 97.92%, and 0.004%, respectively while classified using an ensemble learning method.
V-A3 Hybrid Metaheuristics and ML
@@
In the context of IoT intrusion detection systems, existing hybrid optimization algorithms can be categorized into 6 small groups: swarm+physics-based [75], nature+physics-inspired [45, 17], nature+math-inspired [139, 11], nature+human-based [77, 159], swarm+search-based [28], and nature+search-inspired [10, 20, 129]. Though utilizing physics-based techniques with nature-inspired ones (binary gravitational search (BGSA)+GWO [45], simulated annealing (SA)+shuffled shepherd optimization (SSO) [17]) prove to be effective, integrating with swarm-based optimizers (multi-object PSO+Lévy flight [75]) turns out to be inefficient. Regarding the third group, Rahmani et al. [139] employ grasshopper optimization (GAO) and AOA for parameters and hyperparameters tuning in a random neural network (RdNN), which generates an IDS having 99.56% precision and 99.37% detection rate. In another work [11], a binary chimp optimization algorithm (BCOA) is integrated with the sine cosine algorithm (SCA) for securing the IoT-WSN network. Although the IDS generates high accuracy and specificity (99.63% and 99.67%, respectively), the f1-score is not convincing (94.52%). The well-known human-based metaheuristic object-based learning (OBL) is hybridized with Harris Hawk Optimization (HHO) and Golden Jackal Optimization Algorithm (GJOA), respectively in [159] and [77] to select best possible features. These DT-based and LSTM-driven IoT-IDSs provide remarkable performances with 99.65%100% and 98.93% accuracy, respectively. Baniasadi et al. [28] adjust parameters of deep CNN (DCNN) utilizing neighborhood search (NS)-based PSO. They get a negligible amount of mean square error (0.00053%) with 98.86% accuracy and 95.32% specificity on the UNSW-NB15 and Bot-IoT datasets. Turning to the last group of this category, it can be concluded that selecting features using a combination of nature-inspired and search-based metaheuristics is not effective according to the results provided in [10, 20, 129]. It is worth mentioning that these IDSs use EL classifier, variational autoencoder (VAE), and deep RL, respectively.
@@
Table V shows the list of works in the existing literature that leverage different metaheuristics-based techniques and ML algorithms. According to the table, most systems leverage nature-inspired optimization techniques to increase the efficiency of the classifiers. Besides, swarm-based, population-based hybrid, and hybrid metaheuristics are utilized significantly. Moreover, a thorough analysis of the IoT-IDSs based on different performance metrics, metaheuristics, their applications, ML algorithms, classification types, and datasets are presented in Appendix A (Table LABEL:tab:performance_appendix).
@@
| Swarm-based | [147, 18, 151, 40] | [167, 134, 40] | [166, 161, 89, 136] | [21, 90] | [54, 34] | [54, 23, 29, 146, 12, 43] | |
| Nature-inspired | [69, 112, 7, 149, 78, 55, 156] | [152, 85, 173, 67, 66, 71, 170, 145] | [109, 68, 136, 30, 123, 137, 73] | [20, 128, 37, 124, 56] | [142, 39, 24, 86, 35, 60, 140, 1, 88, 27, 97, 12, 32, 165, 94, 133, 48, 8] | ||
| EA-based | [147, 46] | [73, 106] | [23, 108] | ||||
| Population-based hybrid | [91, 13, 98, 103, 14, 42, 104, 132, 65, 57, 96, 92, 2, 41, 114] | [98, 87] | [144] | [150, 144, 26, 16, 74, 154, 5, 8] | |||
| Phy/Math-based | [63, 115] | [113] | [116] | ||||
| Human-inspired | [131, 93] | [9, 136] | [36] | [59, 62, 29, 36] | |||
| Search-based | [127] | ||||||
| Hybrid | [17, 75, 159] | [45, 11, 10] | [139] | [20] | [129] | [58, 28, 77] | |
@@
We find that most of the existing IoT-IDSs solely utilize different metaheuristics algorithms for selecting an optimal set of features. Besides, some of the systems employ them distinctly for optimizing the parameters and hyperparameters of the machine learning models. However, a few works focus on hybrid applications by leveraging feature selection and parameter (FS-PT) or hyperparameter (FS-HPT) tuning in the same detection system. Figure 6 illustrates the percentage of different applications leveraged by the IDSs, and Table VI shows the list of corresponding existing works.
@@
| FS | [91, 142, 13, 147, 109, 150, 46, 103, 45, 58, 39, 59, 63, 17, 21, 18, 14, 42, 104, 75, 132, 161, 112, 65, 167, 57, 24, 86, 96, 85, 151, 60, 89, 68, 87, 2, 173, 134, 127, 7, 67, 41, 10, 66, 74, 149, 90, 1, 154, 129, 123, 77, 159, 114, 78, 137, 156, 32, 165, 93, 124, 115, 73, 36, 48, 113, 108] |
| PT | [98, 9, 139, 144, 28, 136, 128, 88, 97, 34, 37, 94, 133, 43] |
| HPT | [152, 139, 62, 26, 136, 30, 29, 140, 27, 170, 55, 145, 133, 40, 106] |
| FS and PT | [166, 116, 11, 92, 35, 20, 8] |
| FS and HPT | [54, 69, 11, 16, 131, 23, 146, 71, 5, 12, 56] |
V-B RQ2: What are the most commonly used datasets and evaluation metrics for IoT-IDS assessment?